hockey tickets at a local rink costs 10 dollars for adults and 7 dollars for children under twelve. if 200 tickets were sold with cash receipts of 1550 dollars how many adult tickets were sold?
Found 2 solutions by rothauserc, greenestamps:
Answer by rothauserc(4718)

let x be the number of adult tickets and y be the number of children tickets sold, then
:
1) x + y = 200
:
2) 10x + 7y = 1550
:
solve equation 1 for y
:
y = 200 - x
:
substitute for y in equation 2
:
10x +7(200 - x) = 1550
:
10x +1400 -7x = 1550
:
3x = 150
:
x = 50
:
********************************
there were 50 adult tickets sold
:
check the answer
:
substitute 50 for x and 150 for y in equation 2
:
10*50 + 7*150 = 1550
:
500 + 1050 = 1550
:
1550 = 1550
*********************************
:

Answer by greenestamps(13200)



(1) If all 200 tickets were children's tickets, the total sales would be $1400.
(2) The actual total sales was $1550, which is $150 more than that.
(3) Each adult ticket costs $3 more than each children's ticket.
(4) From (2) and (3), the number of adult tickets sold was 150/3 = 50.